# Multilevel lattice codes from Hurwitz quaternion integers

### v2

@@ II-B Quaternion Algebras
Let $O$ be a maximal order, and let $\mathfrak{P}_{1},...,\mathfrak{P}_{n}\subseteq O$ be distinct prime (two-sided) ideals. Let $\mathfrak{P}=\prod_{i=1}^{n}\mathfrak{P}_{i}^{a_{i}}$, $a_{i}\in\mathbb{Z}$, $i=1,...,n$. If $a_{i}\geq 0$ for all $i=1,...,n$, then there is a ring isomorphism
$O/\mathfrak{P}\cong(O/\mathfrak{P}_{1}^{a_{1}})\times...\times(O/\mathfrak{P}_{n}^{a_{n}}).$
The next result together with the previous Theorem 1 will be used in the proof of Theorem 3 in the next section which extends Construction $\pi_{A}$ to Hurwitz integers.
As it is remarked in [22] given an odd rational prime $p$, there exists an irreducible quaternion $\pi$ such that $Re(\pi)\in\{1,2\}$. This is a consequence of three-square’s Legendre theorem [39].
@@
[…]
@@
It is straightforward to see that this map is linear. Since $p\mathscr{H}$ is a prime ideal and $p\mathscr{H}=(\mathscr{H}\overline{\pi})\pi=(\mathscr{H}\pi)\overline{\pi}$, we have,
@@
$=$ $\Psi(\alpha+p\mathscr{H})+\Psi(\beta+p\mathscr{H})$
Let’s show now that $\Psi$ is surjective. In fact, given $(\alpha+\mathscr{H}\pi,\beta+\mathscr{H}\overline{\pi})$, since $LGCD(\pi,\overline{\pi})=1$, by Bèzout’s theorem, there exists $\gamma_{1},\gamma_{2}\in\mathscr{H}$ such that,
$\gamma_{1}\overline{\pi}+\gamma_{2}\pi=1+p\mathscr{H}$ (7)
If $Re(\pi)=1$ we can take $\gamma_{1}=\gamma_{2}=\gamma=\frac{p+1}{2}\in\mathbb{Z}$ and if $Re(\pi)=2$ we can take $\gamma_{1}=\gamma_{2}=\gamma=\frac{p+3}{4}\in\mathbb{Z}$. Then
$\gamma\overline{\pi}+\gamma\pi=1+p\mathscr{H}$ (*)
@@
$\theta=(\alpha\gamma\overline{\pi}+\beta\gamma\pi)+p\mathscr{H}$
@@
$=$ $(\alpha\gamma\overline{\pi}+\mathscr{H}\pi,\beta\gamma\pi+\mathscr{H}\overline{\pi})$
From (*)(*)( * ), considering the classes in $\mathscr{H}\pi$ and $\mathscr{H}\overline{\pi}$ we have,
$\gamma\overline{\pi}+\mathscr{H}\pi=1+\mathscr{H}\pi$ $\Rightarrow\gamma\overline{\pi}=1+\mathscr{H}\pi$
@@
$=$ $(\alpha+\mathscr{H}\pi,\beta+\mathscr{H}\overline{\pi}),$
To conclude that $\Psi$ is an isomorphism, we use the fact that for any $\alpha\in\mathscr{H}$, the number of classes of $\mathscr{H}/\mathscr{H}\alpha$, denoted by $|\mathscr{H}/\mathscr{H}\alpha|$ is $\text{Nrm}(\alpha)^{2}$ [17]. Therefore, the domain and the image of $\Psi$ have the same number of elements:
@@
$\displaystyle\left|\dfrac{\mathscr{H}}{\mathscr{H}\pi}\right|\cdot\left|\dfrac{\mathscr{H}}{\mathscr{H}\overline{\pi}}\right|=$ $(\text{Nrm}(\pi))^{2}\cdot(\text{Nrm}(\overline{\pi}))^{2}=p^{2}\cdot p^{2}$
@@
Note also that, under the notation in the proof of the last theorem,
@@ III Multilevel lattice codes over Hurwitz integers
Inpired in [1], we propose here an extension of the Construction $\pi_{A}$ to the Hurwitz integers. As seen, Hurwitz integers have good number-theoretical behavior. Therefore, it is natural to investigate this extension considering their applications in communications.
A strong motivation for extending Construction $\pi_{A}$ to Hurwitz integers was also the paper [22]. There it is proposed the construction of lattice codes over $\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$ and $\mathscr{H}$, obtaining the so-called index lattice codes [40, 41]. In [22], it is used lattice constellations along with an injective map which is associated to the Chinese Remainder Theorem (CRT) in the commutative cases.
@@
[…]
$\mathscr{H}/q\mathscr{H}\cong\mathscr{H}/\mathscr{H}\pi_{1}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{1}\times\ldots\times\mathscr{H}/\mathscr{H}\pi_{k}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{k}.$
@@
$\mathscr{H}/q\mathscr{H}$ $\stackrel{{\scriptstyle(a)}}{{\cong}}\mathscr{H}/p_{1}\mathscr{H}\times\dots\times\mathscr{H}/p_{k}\mathscr{H}$
$\stackrel{{\scriptstyle(b)}}{{\cong}}\mathscr{H}/\mathscr{H}\pi_{1}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{1}\times\dots\times\mathscr{H}/\mathscr{H}\pi_{k}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{k},$
where (a) follows from Theorem 1 and (b) from Theorem 2. Therefore, there exists a ring isomorphism $\varphi$ between $\mathscr{H}/q\mathscr{H}$ and the product $\mathscr{H}/\mathscr{H}\pi_{1}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{1}\times\dots\times\mathscr{H}/\mathscr{H}\pi_{k}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{k}$.
@@
Let’s consider distinct odd rational primes $p_{1},\dots,p_{n}$, and let $q=p_{1}\cdot...\cdot p_{n}$, $q\mathscr{H}$ and $p_{i}\mathscr{H}$ be two-sided ideals generated by $q$ and $p_{i}$, respectively, and, let $\mathscr{H}\pi_{i}$ and $\mathscr{H}\overline{\pi}_{i}$ be prime left ideals generated by irreducible elements of $\mathscr{H}$. The steps to obtain an ring isomorphism $\varphi_{b}:\mathscr{H}/q\mathscr{H}\rightarrow\mathscr{H}/\mathscr{H}\pi_{1}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{1}\times\dots\times\mathscr{H}/\mathscr{H}\pi_{k}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{k}$ is as follows.
@@
where $\gamma_{i}\in\mathbb{Z}$, for $i=1,\ldots,k$ and $\gamma_{i}=\frac{p_{i}+1}{2}$ or $\frac{p_{i}+3}{4}$.
An isomorphism regarding Theorem 1 for $O=\mathscr{H}$ can be obtained according to Bézout identity. Since $\text{GCD}(p_{1},p_{2},\ldots,p_{k})=1$, there exist $\zeta_{1},\ldots,\zeta_{n}\in\mathscr{H}$ such that for $\nu_{i}=q/p_{i}$, we have,
@@
Therefore, we can considere the isomorphism
ϕ1::superscriptitalic-$\phi^{-1}:$ $\mathscr{H}/p_{1}\mathscr{H}\times\ldots\times\mathscr{H}/p_{k}\mathscr{H}\rightarrow\mathscr{H}/q\mathscr{H}$
$\phi^{-1}(\hat{a}_{1},\ldots,\hat{a}_{k})$ $=(\nu_{1}\zeta_{1}\hat{a}_{1}+\ldots+\nu_{k}\zeta_{k}\hat{a}_{k})\mod{q\mathscr{H}}.$
Replacing each $\hat{a}_{i}=\Psi^{-1}_{i}(a_{i}^{(1)},a_{i}^{(2)})$, we arrive at the ring isomorphism,
$,\displaystyle\varphi_{b}^{-1}(a_{1}^{(1)},a_{1}^{(2)},$$\ldots,a_{k}^{(1)},a_{k}^{(2)})=$
$=\left[\sum_{i=1}^{k}\nu_{i}\zeta_{i}\gamma_{i}(\overline{\pi}_{i}a_{i}^{(1)}+\pi_{i}a_{i}^{(2)})\right]\mod q\mathscr{H}.$ (9)
Note that $\varphi_{b}$ is the standard modulo map
$\varphi_{b}:\mathscr{H}/q\mathscr{H}$ $\rightarrow\mathscr{H}/\mathscr{H}\pi_{1}\times\ldots\times\mathscr{H}/\mathscr{H}\overline{\pi}_{k}$
$a$ $\mapsto(a\mod\mathscr{H}\pi_{1},\ldots,a\mod\mathscr{H}\overline{\pi}_{k}).$
@@
Let $\mathscr{H}$ be the maximal order of Hurwitz integers, let $p_{1},\ldots,p_{k}$ be distinct odd rational primes and let $q=\prod_{j=1}^{k}p_{j}$. Let $m_{j}$ and $n$ be integers such that $m_{j}^{(i)}\leq n,i=1,2$, and let $G_{j}^{(1)}:n\times m_{j}^{(1)}$ be a generator matrix of a linear code in $(\mathscr{H}/\mathscr{H}\pi_{j})^{n}$ and $G_{j}^{(2)}:n\times m_{j}^{(2)}$ a generator matrix of a linear code in $(\mathscr{H}/\mathscr{H}\overline{\pi}_{j})^{n}$ for $j\in\{1,\ldots,k\}$. Construction $\pi_{A}$ over Hurwitz integers consists of the following steps:
@@
    Define the discrete codebooks $\mathscr{C}^{(1)}_{j}=\{G_{j}^{(1)}\cdot u:u\in(\mathscr{H}/\mathscr{H}\pi_{j})^{m_{j}^{(1)}}\}$ and $\mathscr{C}^{(2)}_{j}=\{G_{j}^{(2)}\cdot u:u\in(\mathscr{H}/\mathscr{H}\overline{\pi}_{j})^{m_{j}^{(2)}}\}$ for $j\in\{1,\ldots,k\}$.
@@
    Construct $\mathscr{C}=\varphi^{-1}(\mathscr{C}^{(1)}_{1},\mathscr{C}^{(2)}_{1},\ldots,\mathscr{C}^{(1)}_{k},\mathscr{C}^{(2)}_{k})\subseteq(\mathscr{H}/q\mathscr{H})^{n}$, where $\varphi:(\mathscr{H}/q\mathscr{H})^{n}\rightarrow(\mathscr{H}/\mathscr{H}\pi_{1})^{n}\times(\mathscr{H}/\mathscr{H}\overline{\pi}_{1})^{n}\times\ldots\times(\mathscr{H}/\mathscr{H}\pi_{k})^{n}\times(\mathscr{H}/\mathscr{H}\overline{\pi}_{k})^{n}$ is a ring isomorphism.
@@
For $q=3$ and $n=2$ consider the following isomorphism,
$\varphi_{b}^{-1}:(\mathscr{H}/\mathscr{H}(1+i+j))^{2}\times(\mathscr{H}/\mathscr{H}(1-i-j))^{2}\rightarrow(\mathscr{H}/3\mathscr{H})^{2}.$
From (8), since $\pi=1+i+j$, with $Re(\pi)=1$, then $\gamma=\left(\frac{3+1}{2}\right)$, and
$\varphi_{b}^{-1}(a^{(1)},a^{(2)})=2\left[(1-i-j)a^{(1)}+(1+i+j)\ a^{(2)}\right]\mod{3\mathscr{H}}.$
where $a^{(1)}=(a_{1}^{(1)},a_{2}^{(1)})\in(\mathscr{H}/\mathscr{H}\pi)^{2}$ and $a^{(2)}=(a_{1}^{(2)},a_{2}^{(2)})\in(\mathscr{H}/\mathscr{H}\overline{\pi})^{2}$. Now, we can use $\varphi_{b}^{-1}$ to obtain a Construction $\pi_{A}$ lattice over Hurwitz integers. Consider $G_{1}=(1,0)^{T}$ and $G_{2}=(1,1)^{T}$ as a generator matrix of linear codes over $(\mathscr{H}/\mathscr{H}\pi)^{2}$ and $(\mathscr{H}/\mathscr{H}\overline{\pi})^{2}$, respectively and the associate codes,
$\mathscr{C}_{1}$ =$=\{(1,0)^{T}\cdot u;u\in\mathscr{H}/\mathscr{H}\pi\}\subseteq(\mathscr{H}/\mathscr{H}\pi)^{2},$
$\mathscr{C}_{2}$ =$=\{(1,1)^{T}\cdot u;u\in\mathscr{H}/\mathscr{H}\overline{\pi}\}\subseteq(\mathscr{H}/\mathscr{H}\overline{\pi})^{2}.$
For $\mathscr{C}=\Psi^{-1}(\mathscr{C}_{1},\mathscr{C}_{2})=\langle(1,-1-i-j)\rangle\subseteq(\mathscr{H}/3\mathscr{H})^{2}$, we have
$\Lambda_{\pi_{A}}^{\mathscr{H}}(\mathscr{C})=\mathscr{C}+3\mathscr{H}^{2},$
which represents the Construction $\pi_{A}$ lattice generated by $\mathscr{C}$. Note that in this case $|\mathscr{C}|=81=3^{4}$ and $|\mathscr{C}_{1}|=|\mathscr{C}_{2}|=3^{2}$.
@@
A construction that utilizes the Lipschitz quaternions natural order was proposed in [5], where it is feasible to construct codes considering $p$ prime such that $p\equiv 1\mod 4$. The construction described in Definition 12 enlarges this one because consider any odd rational prime $p$.

### v3

@@ II-B Quaternion Algebras
Let $O$ be a maximal order, and let $\mathfrak{P}_{1},\ldots,\mathfrak{P}_{n}\subsetneq O$ be distinct (two-sided) prime ideals. Let $\mathfrak{P}=\prod_{i=1}^{n}\mathfrak{P}_{i}^{a_{i}}$, $a_{i}\in\mathbb{Z}$, $i=1,\ldots,n$. If $a_{i}\geq 0$ for all $i=1,\ldots,n$, then there is a ring isomorphism
$O/\mathfrak{P}\cong(O/\mathfrak{P}_{1}^{a_{1}})\times\cdots\times(O/\mathfrak{P}_{n}^{a_{n}}).$
The following result together with the previous Theorem 1 will be used in the proof of Theorem 3 in the next section which extends Construction $\pi_{A}$ to Hurwitz integers.
As it is remarked in [21], given an odd rational prime $p$ there exists an irreducible quaternion $\pi$ with $\text{Nrm}(\pi)=p$ such that $Re(\pi)\in\{1,2\}$. This is a consequence of three-square’s Legendre theorem [40].
@@
where $\mathscr{H}\pi$ and $\mathscr{H}\overline{\pi}$ are left prime ideals with $\text{Nrm}(\pi)=p$ and $Re(\pi)\in\{1,2\}$.
@@
It is straightforward to see that this map is linear. Since $p\mathscr{H}$ is a prime ideal and $p\mathscr{H}=(\mathscr{H}\overline{\pi})\pi=(\mathscr{H}\pi)\overline{\pi}$, we have
@@
$=$ $\Psi(\alpha+p\mathscr{H})+\Psi(\beta+p\mathscr{H}).$
Let’s show now that $\Psi$ is surjective. In fact, given in $(\alpha+\mathscr{H}\pi,\beta+\mathscr{H}\overline{\pi})$, since $LGCD(\pi,\overline{\pi})=1$, by Bèzout’s theorem, there exist $\gamma_{1},\gamma_{2}\in\mathscr{H}$ such that,
$\gamma_{1}\overline{\pi}+\gamma_{2}\pi=1+p\mathscr{H}.$ (7)
If $Re(\pi)=1$, we can take $\gamma_{1}=\gamma_{2}=\gamma=\frac{p+1}{2}\in\mathbb{Z}$ and, if $Re(\pi)=2$ we can take $\gamma_{1}=\gamma_{2}=\gamma=\frac{p+3}{4}\in\mathbb{Z}$. Then
$\gamma\overline{\pi}+\gamma\pi=1+p\mathscr{H},$ (*)
@@
$\theta=(\alpha\gamma\overline{\pi}+\beta\gamma\pi)+p\mathscr{H}.$
@@
$=$ $(\alpha\gamma\overline{\pi}+\mathscr{H}\pi,\beta\gamma\pi+\mathscr{H}\overline{\pi}).$
From ()(*)( ∗ ), considering the classes in $\mathscr{H}\pi$ and $\mathscr{H}\overline{\pi}$ we have
$\gamma\overline{\pi}+\mathscr{H}\pi=1+\mathscr{H}\pi$ $\Rightarrow\gamma\overline{\pi}=1+\mathscr{H}\pi,$
@@
$=$ $(\alpha+\mathscr{H}\pi,\beta+\mathscr{H}\overline{\pi}).$
To conclude that $\Psi$ is an isomorphism, we use the fact that for any $\alpha\in\mathscr{H}$, the number of classes of $\mathscr{H}/\mathscr{H}\alpha$, denoted by $|\mathscr{H}/\mathscr{H}\alpha|$ is $\text{Nrm}(\alpha)^{2}$, [16]. Therefore, the domain and the image of $\Psi$ have the same number of elements,
@@
$\displaystyle\left|\dfrac{\mathscr{H}}{\mathscr{H}\pi}\right|\cdot\left|\dfrac{\mathscr{H}}{\mathscr{H}\overline{\pi}}\right|=$ $(\text{Nrm}(\pi))^{2}\cdot(\text{Nrm}(\overline{\pi}))^{2}=p^{2}\cdot p^{2},$
@@
Note also that, under the notation in the proof of the last theorem for any $(\alpha,\beta)\in\mathscr{H}/\mathscr{H}\pi\times\mathscr{H}/\mathscr{H}\overline{\pi}$,
@@ III Multilevel lattice codes over Hurwitz integers
Inspired in [3], we propose an extension of Construction $\pi_{A}$ to the Hurwitz integers. As seen, the Hurwitz integers have good number-theoretical behaviour. Therefore, it is natural to investigate this extension by considering their applications in communications. Another nice aspect of the Hurwitz integers is their close relationship with the dual lattice $D_{4}^{*}$. In particular, the basis $\{1,i,j,\frac{1+i+j+k}{2}\}$ for $\mathscr{H}$ corresponds to the generator matrix $B_{D_{4}^{*}}=\{(1,0,0,0),(0,1,0,0),(0,0,1,0),(1/2,1/2,1/2,1/2)\}$, [15].
A strong motivation for extending Construction $\pi_{A}$ to Hurwitz integers was also the paper [21]. There, it is proposed the construction of lattice codes over $\mathbb{Z},\mathbb{Z}[i],\mathbb{Z}[\omega]$ and $\mathscr{H}$, obtaining the so-called index lattice codes [41, 42]. In [21], it is used lattice constellations along with an injective map which is associated to the CRT in the commutative cases.
@@
Let $p_{1},p_{2},\ldots,p_{r}$ be distinct odd rational primes and let $q=\prod_{j=1}^{r}p_{j}$. Consider $\pi_{j}\in\mathscr{H}$ such that $\text{Nrm}(\pi_{j})=p_{j}$, for $j=1,\ldots,r$. There exists a ring isomorphism
$\mathscr{H}/q\mathscr{H}\cong\mathscr{H}/\mathscr{H}\pi_{1}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{1}\times\ldots\times\mathscr{H}/\mathscr{H}\pi_{k}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{r}.$
@@
$\mathscr{H}/q\mathscr{H}$ $\stackrel{{\scriptstyle(a)}}{{\cong}}\mathscr{H}/p_{1}\mathscr{H}\times\dots\times\mathscr{H}/p_{r}\mathscr{H}$
$\stackrel{{\scriptstyle(b)}}{{\cong}}\mathscr{H}/\mathscr{H}\pi_{1}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{1}\times\dots\times\mathscr{H}/\mathscr{H}\pi_{k}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{r},$
where (a) follows from Theorem 1 and (b) from Theorem 2. Therefore, there exists a ring isomorphism $\varphi$ between $\mathscr{H}/q\mathscr{H}$ and the product $\mathscr{H}/\mathscr{H}\pi_{1}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{1}\times\dots\times\mathscr{H}/\mathscr{H}\pi_{k}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{r}$.
@@
Let’s consider distinct odd rational primes $p_{1},\dots,p_{r}$, $q=\prod_{i=1}^{r}p_{i}$, $q\mathscr{H}$, and let $p_{i}\mathscr{H}$ be two-sided ideals generated by $q$ and $p_{i}$, respectively. Let $\mathscr{H}\pi_{i}$ and $\mathscr{H}\overline{\pi}_{i}$ be prime left ideals generated by the irreducible elements $\pi_{i},\overline{\pi}_{i}$ of $\mathscr{H}$, $p_{i}=\pi_{i}\cdot\overline{\pi}_{i}$. The steps to obtain an ring isomorphism $\varphi:\mathscr{H}/q\mathscr{H}\rightarrow\mathscr{H}/\mathscr{H}\pi_{1}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{1}\times\dots\times\mathscr{H}/\mathscr{H}\pi_{k}\times\mathscr{H}/\mathscr{H}\overline{\pi}_{r}$ are as follows.
@@
where $\gamma_{i}\in\mathbb{Z}$, for $i=1,\ldots,r$ and $\gamma_{i}=\frac{p_{i}+1}{2}$ or $\frac{p_{i}+3}{4}$.
An isomorphism regarding Theorem 1 for $O=\mathscr{H}$ can be obtained according to Bézout identity. Since $\text{GCD}(p_{1},p_{2},\ldots,p_{r})=1$, there exist $\zeta_{1},\ldots,\zeta_{n}\in\mathscr{H}$ such that for $\nu_{i}=q/p_{i}$, we have
@@
Therefore, we can consider the isomorphism
ϕ1::superscriptitalic-$\phi^{-1}:$ $\mathscr{H}/p_{1}\mathscr{H}\times\ldots\times\mathscr{H}/p_{r}\mathscr{H}\rightarrow\mathscr{H}/q\mathscr{H}$
$\phi^{-1}(\hat{a}_{1},\ldots,\hat{a}_{r})$ $=(\nu_{1}\zeta_{1}\hat{a}_{1}+\ldots+\nu_{k}\zeta_{k}\hat{a}_{r})\mod{q\mathscr{H}}.$
Replacing each $\hat{a}_{i}=\Psi^{-1}_{i}(a_{i}^{(1)},a_{i}^{(2)})$, we arrive at the ring isomorphism
$,\displaystyle\varphi^{-1}(a_{1}^{(1)},a_{1}^{(2)},$$\ldots,a_{k}^{(1)},a_{k}^{(2)})=$
$=\left[\sum_{i=1}^{r}\nu_{i}\zeta_{i}\gamma_{i}(\overline{\pi}_{i}a_{i}^{(1)}+\pi_{i}a_{i}^{(2)})\right]\mod q\mathscr{H}.$ (9)
Note that $\varphi$ is the standard modulo map,
$\varphi:\mathscr{H}/q\mathscr{H}$ $\rightarrow\mathscr{H}/\mathscr{H}\pi_{1}\times\ldots\times\mathscr{H}/\mathscr{H}\overline{\pi}_{r}$
$a$ $\mapsto(a\mod\mathscr{H}\pi_{1},\ldots,a\mod\mathscr{H}\overline{\pi}_{r}).$
@@
Let $\mathscr{H}$ be the maximal order of Hurwitz integers, let $p_{1},\ldots,p_{r}$ be distinct odd rational primes and let $q=\prod_{j=1}^{k}p_{j}$. Let $m_{j}$ and $n$ be integers such that $m_{j}^{(i)}\leq n,i=1,2$, and let $G_{j}^{(1)}:n\times m_{j}^{(1)}$ be a generator matrix of a linear code in $(\mathscr{H}/\mathscr{H}\pi_{j})^{n}$ and $G_{j}^{(2)}:n\times m_{j}^{(2)}$ a generator matrix of a linear code in $(\mathscr{H}/\mathscr{H}\overline{\pi}_{j})^{n}$ for $j\in\{1,\ldots,r\}$. Construction $\pi_{A}$ over Hurwitz integers consists of the following steps:
@@
    Define the discrete codebooks $\mathscr{C}^{(1)}_{j}=\{G_{j}^{(1)}\cdot u:u\in(\mathscr{H}/\mathscr{H}\pi_{j})^{m_{j}^{(1)}}\}$ and $\mathscr{C}^{(2)}_{j}=\{G_{j}^{(2)}\cdot u:u\in(\mathscr{H}/\mathscr{H}\overline{\pi}_{j})^{m_{j}^{(2)}}\}$ for $j\in\{1,\ldots,r\}$.
@@
    Construct $\mathscr{C}=\varphi^{-1}(\mathscr{C}^{(1)}_{1},\mathscr{C}^{(2)}_{1},\ldots,\mathscr{C}^{(1)}_{r},\mathscr{C}^{(2)}_{r})\subset(\mathscr{H}/q\mathscr{H})^{n}$, where $\varphi:(\mathscr{H}/q\mathscr{H})^{n}\rightarrow(\mathscr{H}/\mathscr{H}\pi_{1})^{n}\times(\mathscr{H}/\mathscr{H}\overline{\pi}_{1})^{n}\times\ldots\times(\mathscr{H}/\mathscr{H}\pi_{k})^{n}\times(\mathscr{H}/\mathscr{H}\overline{\pi}_{r})^{n}$ is a ring isomorphism.
@@
Let $q=3$ and $n=1$. Consider the following ring isomorphism,
$\varphi^{-1}:\mathscr{H}/\mathscr{H}(1+i+j)\times\mathscr{H}/\mathscr{H}(1-i-j)\rightarrow\mathscr{H}/3\mathscr{H}.$
From equation (8), since $\pi=1+i+j$ and $\text{Re}(\pi)=1$, we compute $\gamma=\frac{3+1}{2}=2$. Then the inverse isomorphism is given by,
$\varphi^{-1}(a^{(1)},a^{(2)})=2\left[(1-i-j)a^{(1)}+(1+i+j)a^{(2)}\right]\mod{3\mathscr{H}},$
where $a^{(1)}\in\mathscr{H}/\mathscr{H}\pi$ and $a^{(2)}\in\mathscr{H}/\mathscr{H}\overline{\pi}$.
We can now apply $\varphi^{-1}$ to construct a Construction $\pi_{A}$ lattice over the Hurwitz integers. Consider the following linear codes
$\mathscr{C}_{1}$ =$=\left\{\frac{-1-i-j+k}{2}\cdot u\mod\mathscr{H}(1+i+j)\;:\;u\in\mathbb{K}_{9}\right\}$
=$=\left\{0,k,\frac{-1-i-j+k}{2}\right\}\subset\mathscr{H}/\mathscr{H}\pi,$
$\mathscr{C}_{2}$ =$=\left\{\frac{-1-i+j-k}{2}\cdot u\mod\mathscr{H}(1-i-j)\;:\;u\in\mathbb{K}_{9}\right\}$
=$=\left\{0,j,\frac{-1-i+j-k}{2}\right\}\subset\mathscr{H}/\mathscr{H}\overline{\pi}.$
We have $\mathscr{C}=\varphi^{-1}(\mathscr{C}_{1}\times\mathscr{C}_{2})=\{0,-1+i,1-i,-1+j+k,1-j-k,-i+j+k,i-j-k,\frac{1+i+j+k}{2},\frac{1-i-j-k}{2}\}\subset\mathscr{H}/3\mathscr{H}$. Then the associated Construction $\pi_{A}$ lattice is given by:
$\Lambda_{\pi_{A}}^{\mathscr{H}}(\mathscr{C})=\mathscr{C}+3\mathscr{H}^{2}\subset\mathbb{R}^{4},$
which represents the lattice generated by the code $\mathscr{C}$ via Construction $\pi_{A}$. Note that in this example, $|\mathscr{C}|=9=3^{2}$, and $|\mathscr{C}_{1}|=|\mathscr{C}_{2}|=3$.
A geometric visualization of the codes $\mathscr{C}_{1},\mathscr{C}_{2}$ and $\mathscr{C}$, is presented in Figure 1 via a projection onto the imaginary subspace of the quaternions, i.e., onto $\mathbb{R}^{3}$ by omitting the real part of each quaternion.
Figure 1: Projections of the codes of Example 4 onto $\mathbb{R}^{3}$ by omitting the real part. Gray points represent the projection of all 81 elements of $\mathscr{H}/3\mathscr{H}$. In (a), the code $\mathscr{C}_{1}\subset\mathscr{H}/\mathscr{H}\pi$ is shown in blue; in (b), the code $\mathscr{C}_{2}\subset\mathscr{H}/\mathscr{H}\overline{\pi}$ is shown in red, and, in (c) the code $\mathscr{C}=\varphi^{-1}(\mathscr{C}_{1}\times\mathscr{C}_{2})\subset\mathscr{H}/3\mathscr{H}$, obtained via Construction $\pi_{A}$, is represented by black points. The cube serves as a spatial reference in $\mathbb{R}^{3}$.
@@
A construction that exploits the natural order of Lipschitz was proposed in [5]. In that work, space-time index codes were designed for a $2\times 1$ Multiple-Input Single-Output (MISO) channel, based on the Alamout code. There it was considered the principal ideal domain $\mathbb{Z}[i]$ of Gaussian integers, elements $\phi_{1},\ldots,\phi_{r}$ chosen as pairwise coprime, with $q=\prod_{i=1}^{r}\phi_{i}$ and $\text{Nrm}(\phi_{i})=q_{i}$ for $i=1,\ldots,r$ where $q_{i}$ is not necessarily a prime number. Through CRT, it was established an isomorphism,
$\vartheta:\mathbb{Z}[i]/q\mathbb{Z}[i]\cong\mathbb{Z}[i]/\phi_{1}\mathbb{Z}[i]\times\cdots\times\mathbb{Z}[i]/\phi_{r}\mathbb{Z}[i],$
which enables representing, for each $i=1,\ldots,r$, a codeword $w_{i}\in(\mathbb{Z}[i]/\phi_{i}\mathbb{Z}[i])^{2}$ by two components, $w_{i}=(w_{i,0},w_{i,1})$ with $w_{i,l}\in\mathbb{Z}[i]/\phi_{i}\mathbb{Z}[i],l=0,1$. The encoder maps the components $w_{1,l},\ldots,w_{r,l}$ into the signal for the layer $l\in\{0,1\}$ via
$x_{l}=\vartheta(w_{1,l},\ldots,w_{r,l})\in\mathbb{Z}[i]/q\mathbb{Z}[i],$
with the overall codebook being a subset of $\mathscr{C}_{\mathbb{Z}[i]}$ given by
$\mathscr{C}_{\mathbb{Z}[i]}=\left\{\begin{pmatrix}x_{0}&-\overline{x_{1}}\\ x_{1}&\overline{x_{0}}\end{pmatrix};x_{0},x_{1}\in\mathbb{Z}[i]/q\mathbb{Z}[i]\right\}.$
In comparison to this approach, the construction described in Definition 12 does not rely on a layered structure, offering a different approach. This allows the construction of codes using any odd rational prime $p$, enlarging the possibilities compared to the previous framework restricted to Gaussian primes.
Remark 2.
An interesting direction to be considered is the extension of our approach to the maximal order over the octonions $\mathbb{O}$, known as the Cayley integers or Octavian integers [16]. The Octavian integers, denoted by $\mathscr{O}$, are closely related to the notable $E_{8}$ lattice and form a non-commutative, non-associative, but yet alternative [^1] An alternative algebra is a non-associative algebra that must satisfies: $x(xy)=(xx)y$ and $(yx)x=y(xx)$ for all $x$ and $y$ in the algebra [16, p.72]. algebra. This alternative structure preserves enough algebraic properties to enable meaningful arithmetic.
In $\mathscr{O}$, a division algorithm and an ideal theory can still be developed, despite the non-associative nature of the algebra. As shown in [16, pp. 109-110], every left or right ideal in $\mathscr{O}$ is principal, and in fact, all ideals are two-sided. Moreover, any two-sided ideal in $\mathscr{O}$ is of the form $n\mathscr{O}$, where $n$ is a rational integer.
These properties highlight the strong arithmetic structure present in $\mathscr{O}$, enabled by the high symmetry inherited from the $E_{8}$ lattice. Moreover, a version of the CRT can be formulated in this setting, though it mirrors the classical CRT over $\mathbb{Z}$, as remarked in [21].
A significant challenge in extending our construction to this setting lies in characterizing conditions under which a CRT-type factorization, similar to Theorem 2, can be applied to quotient rings of the form $\mathscr{O}/p\mathscr{O}$, where $p=\text{Nrm}(\theta)$ for some Octavian prime $\theta\in\mathscr{O}$. Understanding the arithmetic and algebraic structure of such quotients will be a key step to extends our framework to octonions settings.